OpenStudy (anonymous):

The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

\[\huge y-y_1 = m(x-x_1)\] m is the slope, since we are looking for parallel, they would have the same slope so just plug the points to the equation above

OpenStudy (anonymous):

y + 2 = 2/3(x - 1) ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so now what do I do?

OpenStudy (anonymous):

-2/3 ?

OpenStudy (anonymous):

subtract 2/3 *

OpenStudy (anonymous):

no that's about it, the question didn't specify anything

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

So this is the answer? y + 2 = 2/3(x - 1)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

Do the same thing as before. Make it look like Ax+By=C

OpenStudy (anonymous):

so multiply 2/3 by x - 1 ?

OpenStudy (anonymous):

you'r suppose to put it in a standard form??

OpenStudy (nikato):

Distribut

OpenStudy (anonymous):

wait so what do I do?

OpenStudy (ranga):

y + 2 = 2/3(x - 1) multiply throughout by 3 3y + 6 = 2(x - 1) 3y + 6 = 2x - 2 2x - 3y = 6 + 2 2x - 3y = 8 is the equation in standard form.

OpenStudy (anonymous):

oh I see, thank you!

OpenStudy (anonymous):

And Thank you @nikato and @helpme1.2 Happy New years to you both!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

Happy New Year to You Too

OpenStudy (nikato):

Happy New Years to u guys too!